An isolated conducting sphere whose radius R=1 metre  has charge q=1/ 9 nc ( 1 divide 9) . The energy density at the surface  of sphere is (in joule per metrecube)


(A) ε0/2               (B)ε0


(C)2ε0                (D)ε0/3


 

Grade:

2 Answers

PARTH ARORA
13 years ago

energy density is  given by 1/2 f silen not E^2  where E is the electric field at the surface.

 

E= q^2 /( 4 pief silen not r^2)

 

where r is the radius of the conductor

put values and obtain ans

Abantika Routray
6 years ago
energy density formula =1/2epsilon*E^2                                                                                                                                           1/2*epsilon*(kq/r^2)^2                                                                                                                                                      1/2*epsilon(9*10^9*1/9*10^-9/1)^2
1/2*epsilon.1^2=epsion/2